# CounterCurate: Enhancing Physical and Semantic Visio-Linguistic Compositional Reasoning via Counterfactual Examples

### 20 Feb 2024 v1

@@ Abstract
We propose CounterCurate, a framework to comprehensively improve the visio-linguistic compositional reasoning capability for both contrastive and generative multimodal models. In particular, we identify two under-explored critical problems: the neglect of the physically grounded reasoning (counting and position understanding) and the potential of using highly capable text and image generation models for semantic counterfactual fine-tuning. Our work pioneers an approach that addresses these gaps.
We first spotlight the near-chance performance of multimodal models like CLIP and LLaVA in physically grounded compositional reasoning. We then apply simple data augmentation using a grounded image generation model, GLIGEN, to generate finetuning data, resulting in significant performance improvements: +33% and +37% for CLIP and LLaVA, respectively, on our newly curated Flickr30k-Positions benchmark. Moreover, we exploit the capabilities of high-performing text generation and image generation models, specifically GPT-4V and DALLE-3, to curate challenging semantic counterfactuals, thereby further enhancing compositional reasoning capabilities on benchmarks such as SugarCrepe, where CounterCurate outperforms GPT-4V.
To facilitate future research, we will release our code, dataset, benchmark, and checkpoints at https://github.com/HanSolo9682/CounterCurate.
@@ 1 Introduction
At present, the majority of research focuses on semantic compositional reasoning, leaving another fundamental problem, physically grounded compositional reasoning, under-explored. This includes tasks such as counting and distinguishing left/right and up/down positions between objects. For example, in Figure 1, it is notable that GPT-4V OpenAI (2023c) makes the wrong choice even though it identifies all objects in the image, demonstrating weak physical compositional reasoning.
Figure 2: The data curation pipeline of CounterCurate. Given a positive image-caption pair, we first generate the negative captions, based on which we generate the negative images using the most suitable approach. Specifically, (a) for Flickr30k-Positions (left/right), we flip the positional keyword, and then conduct the horizontal flip for the image; (b) for Flickr30k-Counting, we first manipulate the digit, and then apply grounded image inpainting Li et al. (2023) as the negative image; (c) for Flickr30k-Attributes, we first leverage GPT-4V OpenAI (2023c) to generate reasonable hard negative captions for replacing the noun, replacing the adjective, and swapping the adjectives. Then we leverage DALLE-3 Betker et al. (2023) to generate coherent images.
With such an observation, it is not surprising to find that both contrastive models like CLIP Radford et al. (2021) and generative models like LLaVA Liu et al. (2023b) deliver near random chance performance on our newly curated benchmarks. We hypothesize that modern LMMs are largely oblivious to positional differences. To address this, we curate counterfactual image-text pairs using simple methods such as horizontal flipping for left/right distinctions, and leveraging a grounded image generation model, GLIGEN Li et al. (2023), to accurately replace/remove objects of interest for up/down and counting tasks. Our approach shows significant improvements such as 33% for CLIP and 37% for LLaVA on our Flickr30k-Positions benchmark.
@@
To overcome this, we utilize the high-performance text generation model GPT-4V OpenAI (2023c) and image generation model DALLE-3 Betker et al. (2023) to curate reasonably and sufficiently difficult negatives. Our method empirically demonstrates a significant performance boost by fine-tuning CLIP and LLaVA using our data generation pipeline on benchmarks such as SugarCrepe, where we surpass NegCLIP and GPT-4V.
@@
  • We systematically study physically grounded compositional reasoning, including positional understanding and counting, and highlight the near random performance of current multimodal models, including CLIP and LLaVA, on our newly curated benchmark.
@@
We plan to release our code, dataset, benchmark, and checkpoints upon the paper’s acceptance.
@@ 2.1 Large Multimodal Models
Multimodal models aim at connecting multiple data modalities. For visio-linguistic tasks, multimodal models are constructed either as contrastive Radford et al. (2021); Jia et al. (2021) or generative OpenAI (2023b); Alayrac et al. (2022); Liu et al. (2023b). Recent large multimodal models are trained with large-scale pretraining data and billions of trainable parameters, which significantly improve their generalization capabilities. Contrastive models such as CLIP Radford et al. (2021) connect images and text by aligning the two modality-specific encoders using contrastive loss. Generative models such as LLaVA Liu et al. (2023b, a) embed images into a decoder-only language model as prefix tokens and utilize next-token prediction loss to align the two modalities. Though such models demonstrate impressive capabilities across various vision-language tasks such as retrieval and visual question answering, they perform less desirably over compositional reasoning Diwan et al. (2022); Hsieh et al. (2023).
@@ 2.2 Compositional Reasoning
Compositional reasoning Diwan et al. (2022) in vision and language tasks involves the ability of models to understand and manipulate complex ideas by breaking them down into simpler, constituent components and then recombining them in novel ways. A typical example can be distinguishing “blank pants and white shorts” versus “white pants and blank shorts”. Winoground Diwan et al. (2022) is the pioneering benchmark for compositional reasoning, composed of 400 items, each with two images and two corresponding captions. Given an image, multimodal models are asked to find the matching caption from the provided two options, and vice versa. SugarCrepe Hsieh et al. (2023) and SeeTrue Yarom et al. (2023) further scales the distractive captions by leveraging language models to automatically generate them.
NegCLIP Yuksekgonul et al. (2023) attempts to improve compositional reasoning via fine-tuning CLIP with perturbed captions as the negative text samples in contrastive learning. However, such permuted captions can be easily detected by a text-only transformer Lin et al. (2023), demonstrating that a more effective approach is needed to further enhance compositional reasoning. COCO-Counterfactual Le et al. (2023) further leverages both negative images and negative captions to fine-tune CLIP via a rule-based data curation pipeline.
Moreover, previous works mainly consider composition reasoning with semantically different concepts, rather than physically grounded relationships such as counting, and distinguishing left/right, and above/below. In our paper, we systematically improve both semantic and physical compositional reasoning by leveraging the most capable generative models including GPT-4V OpenAI (2023c) and DALLE-3 Betker et al. (2023).
@@ 2.3 Counterfactual Reasoning
Figure 3: Fine-tuning different types of large multimodal models with CounterCurate. Our pipeline can enhance both contrastive learning models and generative models by augmenting vanilla image-caption data with curated negative images and captions. Specifically, our counterfactual image-caption pairs (a) provide auxiliary contrastive loss for models like CLIP, where positive contrastive units in the similarity matrix are painted as blue/red and negative ones are painted as white, and (b) can be naturally integrated into the original next-token prediction loss in text generations such as LLaVA.
Counterfactual reasoning Morgan and Winship (2015) refers to the process of imagining “what if” scenarios by modifying the input data. In the context of visio-linguistics scenarios, this involves curating negative images and captions by manipulating the original data and observing how it affects the outcome. This helps models understand cause and effect and predict outcomes in situations they’ve never seen before. The key to counterfactual reasoning is to curate meaningful yet hard enough examples. COCO-Counterfactual Le et al. (2023) explores simple linguistic rules to generate negative captions and uses an image editing model Prompt2Prompt Hertz et al. (2022) to produce the negative images Hertz et al. (2022). Prompt2Prompt is less desirable in understanding complex language instructions, making the counterfactual examples less reliable and challenging.
In this paper, we collect the most challenging negative image and text examples by leveraging the most capable language models GPT-4 OpenAI (2023b) and text-to-image generation model DALLE-3 Betker et al. (2023) and GLIGEN Li et al. (2023), which significantly improve our model’s reasoning performance. By utilizing these hard negatives, our model can better grasp the nuances of language and vision, enhancing its performance on tasks that require a sophisticated understanding of the world.
@@ 3 Approach
In Sec. 3.1, we will first introduce the reasoning behind the selected Flickr30k Entities dataset that is used for both benchmarking and fine-tuning. After that, in Sec. 3.2 and Sec. 3.3, we show the near-chance performance of multimodal models like CLIP and LLaVA on physically grounded compositional reasoning tasks, including positional understanding and counting, and show how to improve their performance by generating negative images via data augmentation and grounded image inpainting. Finally, we introduce using capable text and image generation models to improve the semantic compositional reasoning in Sec. 3.4.
@@ 3.1 Necessity of Grounded Image Captions
Previous approaches Yuksekgonul et al. (2023); Hsieh et al. (2023) adopt global image captioned datasets such as COCO-Captions Chen et al. (2015) and craft negative captions to either benchmark or improve compositional reasoning. This method, however, lacks regional information important to physically grounded reasoning, such as determining the relative position of two objects. Therefore, we leverage the grounded image captioning dataset Flickr30k Entities Plummer et al. (2016) to both curate negative image-caption pairs and benchmark models.
@@ 3.2 Improving Positional Understanding
Currently, there is no benchmark that directly evaluates the positional understanding of multimodal models. To build such a benchmark as in Figure 2 (a), we leverage bounding boxes in Flickr30k Entities Plummer et al. (2016) to identify the positional object pairs within each image $I$ that are strictly to the left/right or above/below each other.
@@
Specifically, this process relies on the bounding box coordinates that are formatted as $(x_{j1},y_{j1},x_{j2},y_{j2})$ for every object $j\in I$ whose bounding box’s upper-left and lower-right corner are at $(x_{j1},y_{j1})$ and $(x_{j2},y_{j2})$, respectively. For each pair of objects $(a,b)\in I$, we identify the positional object pairs via the formula:
@@
Based on the formula, we can generate positive and negative positional captions such as “a bike is to the left of a woman” vs “a bike is to the right of a women” and “a table is above a towel” vs “a table is below a towel”.
We also prompt the pure-text GPT4 OpenAI (2023b) model to rewrite the vanilla negative prompt $c_{a}$ is above/below $c_{b}$ to improve the quality of images generated by GLIGEN. For example, the caption “a street is above young child" is very unnatural, while GPT-expanded caption “the street stretches out above a young child, elevated by a bridge” is much better. For detailed prompt engineering, see Appendix B.
@@
Furthermore, we include negative images in our dataset to curate Flickr30k-Positions. To generate negative images for left-right pairs, we apply a simple data augmentation method: flip the original images horizontally. For above-below pairs, however, flipping vertically without context will lead to unreasonable images. To address this issue, we utilize GLIGEN Li et al. (2023), an object-level text-to-image generation model that can ground text prompts with bounding boxes. Specifically, for each above-below object pair $(a,b)\in I$, we obtain their bounding boxes and swap their centers while maintaining the width and height, ensuring that the generated images are more natural, preserving the objects’ sizes and aspect-ratios as demonstrated in Figure 4.
Figure 4: To generate the correct above-below negative image via GLIGEN with prompt "the ball is below the sports outfit", we recenter the bounding boxes of "ball" and "sports outfit" and feed them into GLIGEN together with an expanded caption from GPT4.
@@ 3.3 Improving Object Counting
We introduce the Flickr30k-Counting dataset, which can be used to further enhance the object counting capability for LMMs. Again, we curate counterfactual images and captions as in Figure 2 (b). We count the number of bounding boxes $n_{j}$ for each object category $j\in I$ in the Flickr30k-Entities dataset Plummer et al. (2016). For two distinct object categories $(a,b)$, we generate a positive caption “there are $n_{a}$ $a$ and $n_{b}$ $b$, such as “there are three cats and four dogs” where object $a$ is “cat” and object $b$ is “dog”.
@@
If the two objects do not spatially overlap with any other object in the image, we create our negative caption by decrementing the one with more objects, $b$, and incrementing the one with less objects, $a$; for example, “there are four cats and three dogs”. We only increment/decrement by 1 the number of objects because this enforces a hard counterfactual format where the numbers in the positive captions are reversed in the negatives.
In most cases, however, objects in an image do overlap, and we cannot ensure the increment/decrement rule to generate hard counterfactuals as described above. For example, if a dog overlaps with a cat, and we re-use the method above to remove one dog and add one cat, we might end up only removing one dog along with the overlapping cat unintendedly. In this case, we can only remove one of either $a$ or $b$, for example, the dog $b$, and all objects that it overlaps with. An example of the resulting negative caption is “there are two cats and three dogs.”
@@
In the case where objects do not overlap, we simply leverage GLIGEN Li et al. (2023) to inpaint the category $C_{k}$ where $B_{j}$ was. Otherwise, we remove the target objects by using GLIGEN to replace it with an object from a generic category. In particular, we simply use the plant category for replacement, as plants can appear in almost any natural setting, let it be outdoor or indoor, and we find that GLIGEN does well in generating it according to each image’s background during inpainting. Please refer to Appendix C for more details.
@@ 3.4 Improving Semantic Compositional Reasoning
Existing works Yuksekgonul et al. (2023); Le et al. (2023) use a rule-based approach to generate negative images/captions for semantic compositional reasoning, which results in less desirable negatives such as uncommon negative captions including “A baby with blue hat and orange vest". Here we introduce the Flickr30k-Attributes dataset as in Figure 2 (c), which first utilizes GPT-4V OpenAI (2023c) to generate hard negative captions and then leverages DALLE-3 Betker et al. (2023) for generating corresponding high-quality negative images. We extract the first and most detailed caption $C_{I}$ for each image $I$ from the original Flickr30k dataset Young et al. (2014).
@@
To generate high-quality negative captions, we leverage GPT-4V and feed both positive images and captions to generate negative captions. To let GPT-4V know the objects’ locations, we overlay bounding boxes onto the original positive image Cai et al. (2023) and also feed this annotated image into GPT-4V. We prompt GPT-4V to generate three kinds of captions, including (i) changing a noun, (ii) changing an adjective, and (iii) swapping any of the two phrases, nouns, adjectives, etc. Together, we curate our fine-tuning dataset Flickr30-Attributes. For example, given the original caption “a man wearing a black shirt and blue jeans", GPT-4V can generate corresponding negatives: (i) “a man wearing a black jacket and blue jeans", (ii) “a man wearing a black shirt and red jeans", and (iii) “a man wearing a blue shirt and black jeans". Detailed prompts to GPT-4V are shown in Appendix A.
@@
We simply feed the generated captions from GPT-4V directly into the current most capable text-to-image generation model, DALLE-3 Betker et al. (2023), to generate high-quality images.

### 12 Mar 2024 v2

@@ Abstract
We propose CounterCurate, a framework to comprehensively improve the visio-linguistic compositional reasoning capability for both contrastive and generative multimodal models. In particular, we identify two critical under-explored problems: the neglect of physically grounded reasoning (counting and position understanding) and the potential of using highly capable text and image generation models for semantic counterfactual fine-tuning. Our work pioneers an approach in addressing these gaps.
We first spotlight the near-chance performance of multimodal models like CLIP and LLaVA in physically grounded compositional reasoning. We then apply simple data augmentation using grounded image generation model GLIGEN to generate fine-tuning data, resulting in significant performance improvements: +33% and +37% for CLIP and LLaVA, respectively, on our newly curated Flickr30k-Positions benchmark. Moreover, we exploit the capabilities of high-performing text generation and image generation models, specifically GPT-4V and DALLE-3, to curate challenging semantic counterfactuals, thereby further enhancing compositional reasoning capabilities on benchmarks such as SugarCrepe, where CounterCurate outperforms GPT-4V.
To facilitate future research, we will release our code, dataset, benchmark, and checkpoints at https://countercurate.github.io/.
@@ 1 Introduction
At present, the majority of research focuses on semantic compositional reasoning, leaving another fundamental problem, physically grounded compositional reasoning, under-explored. This includes tasks such as counting and distinguishing left/right and up/down positions between objects. For example, in Figure 1, GPT-4V OpenAI (2023c) notably made the wrong choice even though it identified all objects in question, demonstrating capable semantical yet weak physical compositional reasoning.
Figure 2: The data curation pipeline of CounterCurate. Given a positive image-caption pair, we first generate the negative captions, based on which we curate the negative images using the most suitable approach. Specifically, (a) for Flickr30k-Positions (left/right), we flip the positional keyword before conducting the horizontal flip for the image; (b) for Flickr30k-Counting, we manipulate the digit before applying grounded image inpainting Li et al. (2023) as the negative image; (c) for Flickr30k-Attributes, we first leverage GPT-4V OpenAI (2023c) to generate reasonable hard negative captions for replacing the noun, replacing the adjective, and swapping the adjectives. Then we leverage DALLE-3 Betker et al. (2023) to generate coherent images.
With such an observation, it is not surprising to find that both contrastive models like CLIP Radford et al. (2021) and generative models like LLaVA Liu et al. (2023b) deliver near random chance performance on our newly curated benchmarks. We hypothesize that modern LMMs are largely oblivious to positional differences. To address this, we generate counterfactual image-text pairs using both simple methods, such as horizontal flipping for left/right distinctions, and leveraging a grounded image generation model, GLIGEN Li et al. (2023), to accurately replace/remove objects of interest for up/down and counting tasks. Our approach shows significant improvements such as 33% for CLIP and 37% for LLaVA on our Flickr30k-Positions benchmark.
@@
To overcome this, we utilize high-performance text generation model GPT-4V OpenAI (2023c) and image generation model DALLE-3 Betker et al. (2023) to curate reasonably and sufficiently difficult negatives. Our method empirically demonstrates a significant performance boost by fine-tuning CLIP and LLaVA using our data generation pipeline on benchmarks such as SugarCrepe, where we surpass NegCLIP and GPT-4V.
@@
  • We systematically study physically grounded compositional reasoning, including positional understanding and counting, and highlight the near random performance of multimodal models, including CLIP and LLaVA, on our newly curated benchmarks.
@@ 2.1 Large Multimodal Models
Multimodal models aim at connecting multiple data modalities. For visio-linguistic tasks, multimodal models are constructed either as contrastive Radford et al. (2021); Jia et al. (2021) or generative OpenAI (2023b); Alayrac et al. (2022); Liu et al. (2023b). Recent large multimodal models are trained with large-scale pretraining data and billions of trainable parameters, allowing them to excel at generalization. Contrastive models such as CLIP Radford et al. (2021) connect images and texts by aligning two modality-specific encoders using contrastive loss. Generative models such as LLaVA Liu et al. (2023b, a) embed images into a decoder-only language model as prefix tokens and utilize next-token prediction loss to align the two modalities. Though such models demonstrate impressive capabilities across various vision-language tasks such as retrieval and visual question answering, they perform less desirably over compositional reasoning Diwan et al. (2022); Hsieh et al. (2023).
@@ 2.2 Compositional Reasoning
Compositional reasoning Diwan et al. (2022) in vision and language tasks involves evaluating models’ ability to understand and manipulate complex ideas by breaking them down into simpler, constituent components before recombining them in novel ways. A typical example can be distinguishing “blank pants and white shorts” versus “white pants and blank shorts”. Winoground Diwan et al. (2022) is the pioneering benchmark for compositional reasoning, composed of 400 items, each with two images and two corresponding captions. Given an image, multimodal models are asked to find the matching caption from the provided two options, and vice versa. SugarCrepe Hsieh et al. (2023) and SeeTrue Yarom et al. (2023) further scale the distractive captions by leveraging language models for automated generation.
NegCLIP Yuksekgonul et al. (2023) attempts to improve compositional reasoning via fine-tuning CLIP with perturbed captions as the negative text samples in contrastive learning. However, such permuted captions can be easily detected by a text-only transformer Lin et al. (2023), demonstrating that a more effective approach is needed to further enhance compositional reasoning. DAC Doveh et al. (2024) leverages dense and aligned captions for better positive captions. SGVL Herzig et al. (2023) utilizes scene graph to fine-tune the pretrained vision-language model. TSVLC Doveh et al. (2023) leverages language models to generate positive and negative captions. COCO-Counterfactual Le et al. (2023) further utilizes both negative images and negative captions to fine-tune CLIP via a rule-based data curation pipeline.
Moreover, previous works mainly consider composition reasoning with semantically different concepts rather than physically grounded relationships such as object counting and distinguishing left/right and above/below. Though a recent work Paiss et al. (2023) improved CLIP’s counting capabilities, they neither addressed compositional reasoning nor leveraged negative images. In our paper, we systematically improve both semantic and physical compositional reasoning by leveraging the most capable generative models including GPT-4V OpenAI (2023c) and DALLE-3 Betker et al. (2023).
@@ 2.3 Counterfactual Reasoning
Figure 3: Fine-tuning different types of large multimodal models with CounterCurate. Our pipeline can enhance both contrastive learning models and generative models by augmenting vanilla image-caption pairs with curated negative images and captions. Specifically, our counterfactual image-caption pairs (a) provide auxiliary contrastive loss for models like CLIP, where positive contrastive units in the similarity matrix are painted as blue/red and negative ones are painted as white, and (b) can be naturally integrated into the original next-token prediction loss in text generation models such as LLaVA.
Counterfactual reasoning Morgan and Winship (2015) refers to the process of imagining “what if” scenarios by modifying the input data. In the context of visio-linguistics scenarios, this involves curating negative images and captions by manipulating the original data and observing how it affects the outcome. This helps models understand cause and effect and predict outcomes in situations they’ve never seen before. The key to counterfactual reasoning is curating meaningful and hard enough examples. COCO-Counterfactual Le et al. (2023) explores simple linguistic rules to generate negative captions and uses an image editing model Prompt2Prompt Hertz et al. (2022) to produce negative images Hertz et al. (2022). Prompt2Prompt is less desirable in understanding complex language instructions, making the generated counterfactuals less reliable and challenging.
In this paper, we collect the most challenging negative image and text examples by leveraging the most capable language models GPT-4 OpenAI (2023b), text-to-image generation model DALLE-3 Betker et al. (2023), and GLIGEN Li et al. (2023), significantly improving our model’s reasoning. Utilizing these hard negatives, our model can better grasp the nuances of language and vision, enhancing its performance on tasks that require a sophisticated understanding of the world.
@@ 3 Approach
In Sec. 3.1, we first justify our selecting Flickr30k Entities dataset for both benchmarking and fine-tuning. Further in Sec. 3.2 and 3.3, we demonstrate the near-chance performance of multimodal models like CLIP and LLaVA on physically grounded compositional reasoning tasks, including positional understanding and counting, and improve their performance by generating negative image-text pairs via data augmentation and grounded image inpainting. Finally, we introduce using capable text and image generation models to improve models’ semantic compositional reasoning in Sec. 3.4.
@@ 3.1 Necessity of Grounded Image Captions
Previous approaches Yuksekgonul et al. (2023); Hsieh et al. (2023) adopt global image captioned datasets such as COCO-Captions Chen et al. (2015) and craft negative captions to either benchmark or improve compositional reasoning. This method, however, lacks regional information important to physically grounded reasoning, such as determining the relative position of two objects. Therefore, we leverage Flickr30k Entities Plummer et al. (2016), a grounded image captioning dataset to both curate negative image-caption pairs and benchmark models.
@@ 3.2 Improving Positional Understanding
Currently, there is no benchmark that directly evaluates the positional understanding of multimodal models. We build such a benchmark, Flickr30k-Positions as in Figure 2 (a), that is composed of positive and negative image-caption pairs describing the same objects in opposite positions.
@@
Specifically, we utilize the bounding boxes
$B=\{B_{j}=(x_{j1},y_{j1},x_{j2},y_{j2})\mid\forall j\in I\}$
in Flickr30k Entities Plummer et al. (2016) for each object $j$ in Image $I$ that only has one bounding box, where box $B_{j}$’s upper-left and lower-right corners are at $(x_{j1},y_{j1})$ and $(x_{j2},y_{j2})$, respectively. For each pair of objects $\{a,b\}\subseteq I$, we identify the positional object pairs via the formula:
@@
With the formula, we generate positive and negative positional captions $C$ and $C^{\prime}$ such as $C=$“a bike is to the left of a woman” vs. $C^{\prime}=$“a bike is to the right of a women” and $C=$“a table is above a towel” vs. $C^{\prime}=$“a table is below a towel”.
We also prompt text-only GPT4 OpenAI (2023b) to rewrite the vanilla negative prompt $a$ is above/below $b$ to better guide image generation models. For example, the caption “a street is above young child" is very unnatural, while GPT-expanded caption “the street stretches out above a young child, elevated by a bridge” is much better. For detailed prompt engineering, see Appendix B.
@@
Furthermore, we include negative images in our dataset to curate Flickr30k-Positions. For the left-right pairs, we apply a simple data augmentation method: flip the original images horizontally. For the above-below pairs, however, flipping vertically without context would lead to unreasonable images. To address this issue, we utilize GLIGEN Li et al. (2023), an object-level text-to-image generation model that can ground text prompts with bounding boxes. In particular, for each above-below object pair $(a,b)\in I$, we swap the centers of the bounding boxes $B_{a},B_{b}$ while maintaining their widths and heights, ensuring that the generated images are natural while preserving the objects’ sizes and aspect-ratios as demonstrated in Figure 4. The new bounding boxes and prompts are fed to GLIGEN to generate the corresponding negative image $I^{\prime}$.
Figure 4: To generate the correct above-below negative image via GLIGEN with the original prompt "the ball is below the sports outfit", we recenter the bounding boxes of "ball" and "sports outfit" and feed them into GLIGEN together with an expanded prompt from GPT4.
@@ 3.3 Improving Object Counting
We introduce Flickr30k-Counting to enhance LMMs’ object counting capabilities. We again curate counterfactual image-caption pairs as in Figure 2 (b). We count the number of bounding boxes for each object category in Flickr30k-Entities Plummer et al. (2016). For two distinct object categories $\{P,Q\}\subseteq I$, we generate a positive caption $C=$“there are $n_{P}$ $P$’s and $n_{Q}$ $Q$’s”, such as “there are three cats and four dogs” where category $P$ is “cat” and category $Q$ is “dog”.
@@
If both objects do not spatially overlap with other objects, we generate the corresponding negative caption by decrementing the one with more objects, $Q$, and incrementing the one with fewer, $P$; for example, $C^{\prime}=$“there are four cats and three dogs”. This enforces a hard counterfactual format as the numbers in the positives are reversed in the negatives.
In most cases, however, objects in an image do overlap, and we cannot ascertain that the increment/decrement rule still generates hard counterfactuals. For example, if a dog overlaps with a cat, and we re-use the rule above to remove one dog and add one cat, we end up with “three cats and three dogs”. In this case, we can only remove either one of $P$ or one of $Q$, for example, the dog, and all objects it overlaps with. An example of the resulting negative caption is $C^{\prime}=$“there are two cats and three dogs.”
@@
In the case where objects do not overlap, we simply leverage GLIGEN Li et al. (2023) to inpaint a new object of the incremented category $P$ at the bounding box of one of the objects from the decremented category $Q$. Otherwise, we remove the target objects using GLIGEN to replace it with an object from a generic category. We simply use the plant category for replacement, as plants can appear in almost any natural setting, let it be outdoor or indoor, and we find that GLIGEN does well in generating it according to each image $I$’s background during inpainting for each corresponding negative image $I^{\prime}$. More details in Appendix C.
@@ 3.4 Improving Semantic Compositional Reasoning
Existing works Yuksekgonul et al. (2023); Le et al. (2023) use a rule-based approach to generate negative images/captions for semantic compositional reasoning, resulting in less desirable negatives such as uncommon negative captions including “A dog with blue fur" from “A dog with black fur". We introduce Flickr30k-Attributes as in Figure 2 (c), which utilizes GPT-4V OpenAI (2023c) to generate hard negative captions and leverages DALLE-3 Betker et al. (2023) for generating corresponding high-quality negative images. We extract the first and most detailed caption $C$ for each image $I$ from the original Flickr30k dataset Young et al. (2014).
@@
To generate high-quality negative captions, we leverage GPT-4V and feed both positive images $I$ and captions $C$ to generate negative captions $C^{\prime}$. To let GPT-4V know the objects’ locations, we overlay bounding boxes $B$ onto the original positive image Cai et al. (2023) and also feed this annotated image into GPT-4V. We prompt GPT-4V to generate three kinds of captions, including (i) changing a noun, (ii) changing an adjective, and (iii) swapping any of the two phrases, nouns, adjectives, etc. For example, given the original caption $C$ “a man wearing a black shirt and blue jeans", GPT-4V can generate corresponding negatives $C^{\prime}$: (i) “a man wearing a black jacket and blue jeans", (ii) “a man wearing a black shirt and red jeans", and (iii) “a man wearing a blue shirt and black jeans". Detailed prompts to GPT-4V are shown in Appendix A.
@@
We simply feed the curated captions $C^{\prime}$ from GPT-4V directly into the currently most capable text-to-image generation model, DALLE-3 Betker et al. (2023), to generate high-quality negative images $I^{\prime}$.